# Omni-Router: Sharing Routing Decisions in Sparse Mixture-of-Experts for Speech Recognition

### v2

@@ V-D Evaluation
VI Results
VI-A Omni-router improves ASR
In Table I, we present the WER results comparing a baseline dense ASR model (140M) with Switch Transformer and Omni-router Transformer ASR models, each employing two experts per layer. We first note that both MoE-based architectures consistently outperform the dense baseline across all evaluated out-of-domain datasets. Additionally, the Omni-router Transformer further improves upon the performance of the Switch Transformer. Given that the Omni-router shares hyperparameters originally optimized for the Switch Transformer, its superior performance despite minimal hyperparameter tuning underscores the effectiveness and robustness of the shared-routing strategy.
VI-B Why Does the Omni-router Transformer Outperform?
A central question raised by our experimental results is what factors contribute to the superior performance of the Omni-router Transformer. We hypothesize that the shared routing mechanism creates stable and structured ”pathways” for tokens through the network, enabling the model to develop more cooperative and specialized experts.
@@ VI-B Why Does the Omni-router Transformer Outperform?
Our visualizations reveal significant differences in routing behavior between the two architectures. The Omni-router Transformer exhibits notably structured and coherent expert assignments across both temporal and depth dimensions, whereas the Switch Transformer demonstrates a more fragmented and less organized routing pattern. Specifically, the Omni-router’s routing creates consistent temporal segments, suggesting meaningful coordination and communication between layers. Furthermore, repeated utilization of the same experts across multiple layers indicates that the Omni-router’s inductive bias effectively promotes stable specialization. Crucially, expert assignments in the Omni-router Transformer clearly correlate with acoustic patterns in the mel-spectrogram. For example, the ”green” expert predominantly handles silent segments, while alternating patterns of orange, blue, and pink experts correspond clearly to distinct speech segments. In contrast, the Switch Transformer displays fragmented expert assignments, indicating weaker inter-layer coordination and suboptimal specialization.
@@
Greater Expert Specialization: To rigorously quantify expert specialization, we adopted an expert permutation method proposed in prior research [15]. Specifically, we randomly permuted expert assignments for each token with varying probability $p$ and measured the resulting impact on ASR performance, quantified by WER degradation. We hypothesized that the Omni-router Transformer would exhibit a greater sensitivity to incorrect expert assignments, reflecting higher expert specialization.
Figure 6: Relative WER change percentage of Switch Transformer ASR and Omni-router Transformer ASR at different permutation ratios $p$.
@@
VI-C Omni-router Transformer outperforms Switch Transformer across different number of experts
In Table II, we investigate how varying the number of experts impacts the performance of Switch Transformer and Omni-router Transformer-based ASR models, compared against a baseline dense ASR model (84M parameters). We present results on the Librispeech dataset (we observed similar trends across other datasets). Our findings demonstrate that Omni-router ASR consistently achieves superior performance over Switch Transformer ASR across all evaluated expert configurations. Notably, Switch Transformer’s performance deteriorates at 8 experts, while Omni-router’s structured layer-coupled routing effectively mitigates this degradation, maintaining stable performance even when using more experts.
TABLE II: WERs of Switch Transformer ASR and Omni-router Transformer ASR models with different number of experts.
@@ VI-C Omni-router Transformer outperforms Switch Transformer across different number of experts
VI-D Omni-router Transformer outperforms Switch Transformer across different model sizes
In Table III, we examine the impact of model size by evaluating baseline dense ASR models with parameter counts of 84M, 140M, and 200M. Correspondingly, we train sets of 2-expert Switch Transformer and 2-expert Omni-router Transformer ASR models matched to these sizes. We find that the Omni-router ASR models consistently achieve lower WERs compared to both the Switch Transformer ASR models and the baseline dense ASR models.
TABLE III: WERs of Dense models, and Switch Transformer and Omni-router Transformer ASR models with different number of model sizes and two experts.
@@ VI-D Omni-router Transformer outperforms Switch Transformer across different model sizes
VI-E Results are reproducible on single and specialized datasets
To validate the generalizability and robustness of the Omni-router Transformer beyond conversational data, we extended our evaluation to the domain of read speech. For this, we conducted experiments on the Libriheavy dataset [52], a large-scale ASR corpus derived from audiobooks in the LibriVox project  [^1] https://librivox.org. We trained three models, each with approximately 84M active parameters: a dense baseline, an 8-expert Switch Transformer, and an 8-expert Omni-router Transformer. The models were trained on the Libriheavy training set, and evaluated on Librispeech dev and test splits.
TABLE IV: WERs of a dense ASR model, a Switch Transformer ASR model and a Omni-router Transformer ASR model trained on Libriheavy and evaluated on Librispeech dataset.
@@ VI-E Results are reproducible on single and specialized datasets
As shown in Table IV, the Omni-router Transformer consistently achieves the lowest WER across all evaluation sets. Particularly, on dev/test-other splits, Omni-router Transformer achieves a relative WER reduction of 13.1%/8.8% over the Switch Transformer model and 19.7%/14.7% over the dense baseline. These results confirm that Omni-router Transformer’s routing mechanism can successfully translate to the distinct characteristics of read speech, which is a more robust and effective architecture for ASR tasks in diverse acoustic environments.
VI-F Omni-router Transformer demonstrates better robustness
@@
To investigate these possibilities, we first relabeled Librispeech data using WhisperX following the same procedure applied to SpeechCrawl, subsequently training Switch Transformer ASR models. The resulting training loss remained stable, suggesting that the pseudo labels alone are unlikely to be the primary cause of instability. Next, we trained the model using real conversational datasets such as Fisher [53], observing training behaviors similar to those encountered with SpeechCrawl, including the characteristic early-stage spike. These findings indicate that data characteristics, rather than label quality, likely drive the observed instability. Conversely, Omni-router Transformer models consistently demonstrated stable training behavior across various data types, highlighting their robustness in large-scale ASR training scenarios.
VII Comparisons with Dense Models
@@
Dataset Omni-router (246M - 140M) Dense (210M) Whisper-small.en (244M)
@@
Librispeech (clean) 3.3 3.7 3.1
Librispeech (other) 7.3 8 7.4
@@
Tedlium 4.2 4.3 4
Voxpopuli 8.5 9 8.2
VIII Conclusion
@@ Acknowledgment
The authors thank Dan Busbridge, Jason Ramapuram and Russ Webb for their insightful comments and Shuangfei Zhai for their assistance with the internal publication process.

### v3

@@ 6.1 Omni-router improves ASR
6.2 Why Does the Omni-router Transformer Outperform
A central question raised by our experimental results is what factors contribute to the superior performance of the Omni-router Transformer. We hypothesize that the shared routing mechanism creates stable and structured "pathways" for tokens through the network, enabling the model to develop more cooperative and specialized experts.
@@ 6.2 Why Does the Omni-router Transformer Outperform
Our visualizations reveal significant differences in routing behavior between the two architectures. The Omni-router Transformer exhibits notably structured and coherent expert assignments across both temporal and depth dimensions, whereas the Switch Transformer demonstrates a more fragmented and less organized routing pattern. Specifically, the Omni-router’s routing creates consistent temporal segments, suggesting meaningful coordination and communication between layers. Furthermore, repeated utilization of the same experts across multiple layers indicates that the Omni-router’s inductive bias effectively promotes stable specialization. Crucially, expert assignments in the Omni-router Transformer clearly correlate with acoustic patterns in the mel-spectrogram. For example, the "green" expert predominantly handles silent segments, while alternating patterns of orange, blue, and pink experts correspond clearly to distinct speech segments. In contrast, the Switch Transformer displays fragmented expert assignments, indicating weaker inter-layer coordination and suboptimal specialization.
@@
Greater Expert Specialization: To rigorously quantify expert specialization, we adopted an expert permutation method proposed in prior research (Dai et al.,, 2024). Specifically, we randomly permuted expert assignments for each token with varying probability $p$ and measured the resulting impact on ASR performance, quantified by WER degradation. We hypothesized that the Omni-router Transformer would exhibit a greater sensitivity to incorrect expert assignments, reflecting higher expert specialization.
Figure 6: Relative WER change percentage of Switch Transformer ASR and Omni-router Transformer ASR at different permutation ratios pp.
@@
6.3 Omni-router Transformer outperforms Switch Transformer across different number of experts
In Table 2, we investigate how varying the number of experts impacts the performance of Switch Transformer and Omni-router Transformer-based ASR models, compared against a baseline dense ASR model (84M parameters). We present results on the Librispeech dataset (we observed similar trends across other datasets). Our findings demonstrate that Omni-router ASR consistently achieves superior performance over Switch Transformer ASR across all evaluated expert configurations. Notably, Switch Transformer’s performance deteriorates at 8 experts, while Omni-router’s structured layer-coupled routing effectively mitigates this degradation, maintaining stable performance even when using more experts.
Table 2: WERs of Switch Transformer ASR and Omni-router Transformer ASR models with different number of experts.
@@ 6.3 Omni-router Transformer outperforms Switch Transformer across different number of experts
6.4 Omni-router Transformer outperforms Switch Transformer across different model sizes
In Table 3, we examine the impact of model size by evaluating baseline dense ASR models with parameter counts of 84M, 140M, and 200M. Correspondingly, we train sets of 2-expert Switch Transformer and 2-expert Omni-router Transformer ASR models matched to these sizes. We find that the Omni-router ASR models consistently achieve lower WERs compared to both the Switch Transformer ASR models and the baseline dense ASR models.
Table 3: WERs of Dense models, and Switch Transformer and Omni-router Transformer ASR models with different number of model sizes and two experts.
@@ 6.4 Omni-router Transformer outperforms Switch Transformer across different model sizes
6.5 Results are reproducible on single and specialized datasets
To validate the generalizability and robustness of the Omni-router Transformer beyond conversational data, we extended our evaluation to the domain of read speech. For this, we conducted experiments on the Libriheavy dataset (Kang et al.,, 2024), a large-scale ASR corpus derived from audiobooks in the LibriVox project  [^1] https://librivox.org. We trained three models, each with approximately 84M active parameters: a dense baseline, an 8-expert Switch Transformer, and an 8-expert Omni-router Transformer. The models were trained on the Libriheavy training set, and evaluated on Librispeech dev and test splits.
Table 4: WERs of a dense ASR model, a Switch Transformer ASR model and a Omni-router Transformer ASR model trained on Libriheavy and evaluated on Librispeech dataset.
@@ 6.5 Results are reproducible on single and specialized datasets
As shown in Table 4, the Omni-router Transformer consistently achieves the lowest WER across all evaluation sets. Particularly, on dev/test-other splits, Omni-router Transformer achieves a relative WER reduction of 13.1%/8.8% over the Switch Transformer model and 19.7%/14.7% over the dense baseline. These results confirm that Omni-router Transformer’s routing mechanism can successfully translate to the distinct characteristics of read speech, which is a more robust and effective architecture for ASR tasks in diverse acoustic environments.
6.6 Omni-router Transformer demonstrates better robustness
@@
To investigate these possibilities, we first relabeled Librispeech data using WhisperX following the same procedure applied to SpeechCrawl, subsequently training Switch Transformer ASR models. The resulting training loss remained stable, suggesting that the pseudo labels alone are unlikely to be the primary cause of instability. Next, we trained the model using real conversational datasets such as Fisher (Cieri et al.,, 2004), observing training behaviors similar to those encountered with SpeechCrawl, including the characteristic early-stage spike. These findings indicate that data characteristics, rather than label quality, likely drive the observed instability. Conversely, Omni-router Transformer models consistently demonstrated stable training behavior across various data types, highlighting their robustness in large-scale ASR training scenarios.
7 Comparisons with Dense Models
@@
Table 5: WER comparisons between a sparse Omni-router ASR (246M total parameters with 140M active parameters) and dense models (210M and 244M).
Dataset
Omni-router
(246M - 140M)
Dense
Whisper-small.en
@@
Librispeech
3.3 3.7 3.1
Librispeech
7.3 8.0 7.4
@@
Tedlium 4.2 4.3 4.0
Voxpopuli 8.5 9.0 8.2
8 Conclusion
@@ Acknowledgment
The authors thank Dan Busbridge, Jason Ramapuram, Russ Webb, and Shuangfei Zhai for their feedback.